# MAVias: Mitigate any Visual Bias

### v1

@@ 3.2.2 Bias Mitigation
$\mathcal{L}^{(i)}_{\text{reg}}=\frac{1}{2}\left\|\|\mathbf{z}^{(i)}_{\text{main}}\|-\lambda\cdot\|\mathbf{z}^{(i)}_{\text{tag}}\|\right\|^{2}$ (3)
@@
4 Experiments
4.1 Datasets and Evaluation Protocol
We employ Biased-CelebA [16], Waterbirds [37], UrbanCars [24], and Imagenet9 [47]. Biased-CelebA is a subset of the well-known CelebA dataset consisting of facial images annotated with 40 binary attributes. In this subset, gender is treated as the target attribute, while the BlondHair attribute introduces bias by enforcing a 90% correlation between the two. Waterbirds is a dataset that features a co-occurrence of 95% between waterbirds (or landbirds) and aquatic environments (or terrestrial environments) as background. UrbanCars is an artificially generated dataset featuring a 95% co-occurrence between car body types and background and/or certain objects relevant to urban or rural regions. ImageNet9 consists of 9 coarse ImageNet classes (dog, bird, vehicle, reptile, carnivore, insect, instrument, primate, and fish).
@@ 4.1 Datasets and Evaluation Protocol
Regarding the evaluation, we follow the protocols suggested by the dataset providers and previous works. In particular, for CelebA we use the accuracy of the underrepresented groups (i.e., bias-conflicting accuracy) and the average accuracy across all groups (i.e., unbiased accuracy) [39, 16]. For Waterbirds, we use the Worst-Group (WG) accuracy and the average accuracy across all groups. In the case of UrbanCars, we calculate the weighted average accuracy across groups, referred to as In-Distribution Accuracy (I.D. Acc), with weights determined by group representation ratios. Furthermore, I.D. Acc serves as a baseline for assessing accuracy drops related to background (BG Gap), co-occurring objects (CoObj Gap), and both background and co-occurring objects combined (BG+CoObj Gap). Lastly, for ImageNet9, we use 7 test set variations for evaluating the impact of background on model’s predictions: ORIGINAL, ONLY-BG-B (black object’s bounding box), ONLY-BG-T (impainted object’s bounding box), NO-FG Black (removed segmented object), ONLY-FG (black background), MIXED-RAND (random background of a random class), and MIXED-NEXT (random background of the next class).
@@ 4.2 Comparative analysis
In this subsection, we compare the performance of MAVias with other competitive BLU methods. Comprehensive results, including comparisons with BLA methods, are provided in the Appendix. Table 2 presents the performance of MAVias on CelebA with gender as target and the BlondHair as the biased attribute. The performance on unbiased and bias-conflicting samples indicates the ability of MAVias to reduce reliance on spurious correlations between the target and the corresponding biased attributes. Specifically, MAVias achieves 87.1% accuracy on bias-conflicting samples, surpassing state-of-the-art BLU methods such as FLAC-B [39] and LfF [31], while maintaining high unbiased accuracy (89.7%).
Table 2: Evaluation on CelebA with BlondHair and gender as biased attribute and target, respectively.
Methods Unbiased Bias-conflict
LfF [31] 84.2$\pm$0.3 81.2$\pm$1.4
SoftCon [16] 84.1 74.4
FLAC-B [39] 87.0$\pm$0.6 84.9$\pm$2.2
MAVias 89.7$\pm$0.6 87.1$\pm$1.7
Table 3: Evaluation on Waterbirds.
Method WG Acc. Avg. Acc.
JTT [26] 86.7$\pm$1.5 93.3$\pm$0.3
DISC [46] 88.7$\pm$0.4 93.8$\pm$0.7
MAVias 93.7$\pm$0.4 94.5$\pm$0.4
Table 3 evaluates MAVias on Waterbirds, a well-known benchmark for spurious correlation problems between bird type and background environment. Notably, MAVias attains 93.7% Worst-Group accuracy, outperforming all compared BLU methods (+5%), while slightly improving the average accuracy (+0.7%). The results on UrbanCars in Tab. 4 further validate the efficacy of MAVias in mitigating multiple biases introduced by co-occurring objects and background features. While the in-distribution accuracy (I.D. Acc) of MAVias is marginally lower than that of LfF [31] or Debian [23], it achieves substantial improvements in reducing the BG Gap (-4.1) and CoObj Gap (-2.4), as well as their combined effects (BG+CoObj Gap: -6.7).
Table 4: Evaluation on UrbanCars.
Method I.D. Acc BG Gap CoObj Gap BG+CoObj Gap
LfF [31] 97.2 -11.6 -18.4 -63.2
JTT [26] 95.9 -8.1 -13.3 -40.1
Debian [23] 98.0 -14.9 -10.5 -69.0
MAVias 92.8$\pm$0.8 -4.1$\pm$0.6 -2.4$\pm$1.4 -6.7$\pm$1.4
Table 5: Accuracy on ImageNet9 test sets. The reported results for the compared methods are based on reimplementations using code provided by the authors of [24].
Vanilla 82.66 $\pm$0.1 85.06 $\pm$0.0 64.16 $\pm$0.1 35.18 $\pm$0.1 44.74 $\pm$0.1 93.12 $\pm$0.0 97.69$\pm$0.0
LfF [31] 78.70 $\pm$0.1 81.47 $\pm$0.2 61.07 $\pm$0.1 34.82 $\pm$0.2 44.46 $\pm$0.0 88.99 $\pm$0.2 94.34 $\pm$0.2
JTT [26] 84.43 $\pm$0.1 86.16 $\pm$0.5 61.09 $\pm$2.0 32.04 $\pm$1.0 36.62 $\pm$4.7 92.09 $\pm$0.5 97.71 $\pm$0.1
Debian [23] 83.02 $\pm$0.4 85.64 $\pm$0.3 64.53$\pm$0.4 34.45$\pm$0.1 45.00 $\pm$0.6 93.06 $\pm$0.1 97.89 $\pm$0.1
MAVias 88.26$\pm$0.1 (+5.24) 89.64$\pm$0.2 (+3.48) 53.02$\pm$0.7 (+8.05) 21.83$\pm$0.4 (-10.21) 32.48$\pm$0.6 (-4.14) 91.90$\pm$0.4 (-1.22) 96.92$\pm$0.2 (-0.97)
@@
In addition, we examine the logit space of the models to demonstrate the effect of MAVias on the behavior of the main model. Similar to the previous experiment, the urban car class of the UrbanCars dataset is used. As one may observe in Fig. 3 (a), the vanilla (biased) model increases the logit values proportionally with the bias, ranging from  4 to  7. This is an expected behavior, as the biases (i.e., spurious correlations) are easier to learn than the actual target and thus the model illustrates higher confidence (or logit values) for the bias-aligned samples. This phenomenon is referred to as Gradient Starvation [34] and can describe how bias is introduced to the models. Also, it is theoretically shown that methodologies penalizing high logit values can mitigate this phenomenon [34]. As shown in Fig. 3 (b), the interactions between the main model and the projection layer in the training procedure (defined by Eq. s (2) and (3)) successfully reduce the main model’s logits values proportionally to the bias in the data.
Apart from the typical training datasets designed for BLA methods, we also consider experiments on ImageNet9, a widely used CV dataset known for various tasks and equipped with multiple test sets specifically designed for bias evaluation. As shown in Tab. 5, MAVias achieves great improvements in the model’s accuracy on the test sets with modified backgrounds (i.e., MIXED-NEXT and MIXED-RAND), with the highest gains occurring for the MIXED-NEXT set (+5.24%) compared to the second best performing method. Similarly, for the sets where background information is suppressed (ONLY-BG-B, ONLY-BG-T, and NO-FG) MAVias achieves improvements of +10.21%, +4.14%, and +8.05%, respectively. As observed, all competitive methods, including MAVias, demonstrate reduced performance on ONLY-FG compared to the vanilla model. This is likely because the vanilla model can exploit biases present in the foreground (e.g., colors) to increase its accuracy. Finally, there is a 0.97% drop in performance on the ORIGINAL test set, which aligns with expectations, as MAVias is designed to rely less on shortcuts that boost overall performance. Furthermore, we evaluate MAVias using two subsets of ImageNet9 ORIGINAL, constructed based on tags identified by MAVias. Specifically, the bias-aligned set includes samples with tags for which the vanilla model performs above its overall accuracy, suggesting high bias impact; the rest of the samples form the bias-conflicting set. Table 6 demonstrates that MAVias improves accuracy on bias-conflicting samples by reducing reliance on visual biases, with a slight decrease on bias-aligned samples for the same reason. Finally, to gain a clearer understanding, Fig. 4 visualizes the bias-aligned tags for the Dog class in the bias-aligned set.
Table 6: Evaluation on the ImageNet9 ORIGINAL set comparing accuracy on the samples with irrelevant tags for which the Vanilla model overperforms (bias-aligned, BA) and the rest of the samples (bias-conflicting, BC).
Method BA acc. BC acc. Gap
Vanilla 98.55 $\pm$0.04 90.86 $\pm$0.11 -7.69 $\pm$0.16
MAVias 97.3 $\pm$0.21 94.13 $\pm$0.01 -3.17 $\pm$0.21
Figure 4: Visualization of the biases identified using MAVias for the Dog class in the ImageNet9 dataset.
@@ 4.3 Ablation Analysis
An important aspect in MAVias is how the embeddings are calculated. Table 7 compares two well-known vison-language models, i.e., CLIP and SigLip, and the text-based model BERT.
@@
BERT [7] 91.7$\pm$1.3 92.8$\pm$1.0
SigLip [48] 92.7$\pm$1.2 93.5$\pm$1.0
CLIP [36] 93.7$\pm$0.4 94.5$\pm$0.4
Furthermore, Tab. 8 compares MAVias’s performance using various LLMs. ChatGPT-4o shows the highest performance, with open-source models also performing competitively. For experiments, we use ChatGPT-4o to minimize LLM-related errors.
@@
Qwen2.5 [17] 7b 91.9$\pm$2.4 94.4$\pm$ 0.5
Mistral-small [18] 22b 92.9$\pm$0.4 94.9$\pm$0.2
Gemma2 [43] 9b 93.5$\pm$0.6 94.5$\pm$0.6
Llama3.1 [10] 8b 93.6$\pm$0.6 94.4$\pm$0.4
ChatGPT-4o [33] >>much-greater-than$>$ >175b 93.7$\pm$0.4 94.5$\pm$0.4
Finally, as discussed in Sec. 3, the regularization term in Eq.  (3) is crucial for balancing the logits between the two models. Table 9 shows the impact of removing this term from the training on the effectiveness of MAVias.

### v2

@@ 3.2.2 Bias Mitigation
$\mathcal{L}^{(i)}_{align}=\frac{1}{2}\left\|\|\mathbf{z}^{(i)}_{\text{main}}\|-\lambda\cdot\|\mathbf{z}^{(i)}_{\text{tag}}\|\right\|^{2}$ (2)
where $\|\cdot\|$ denotes the $l^{2}$-norm, $\lambda\in(0,1)$ is a scaling factor and $\alpha\in(0,1)$ is a weighting factor that balances the influence of the $\mathcal{L}_{align}(\mathbf{z}^{(i)}_{\text{main}},\mathbf{z}^{(i)}_{\text{tag}})$ in the total loss function. Typically, the greater the bias in the data, the smaller the value of $\lambda$ should be, as this leads to smaller $\mathbf{z}_{main}$ values and reduced gradient updates for the bias-aligned samples. Details on the effects of hyperparameters $\lambda$ and $\alpha$ are provided in the supp. material. The overview of the proposed framework is illustrated in Fig. 2.
4 Experiments
4.1 Datasets
We use Biased-CelebA [hong2021bb], Waterbirds [sagawa2019distributionally], UrbanCars [li2023whac], and Imagenet9 [xiao2021noise]. Biased-CelebA is a subset of the CelebA dataset, containing facial images annotated with 40 binary attributes. In this subset, BlondHair is the target attribute, while the gender attribute introduces bias with a 90% correlation. Waterbirds features a 95% co-occurrence between waterbirds (or landbirds) and aquatic (or terrestrial) backgrounds. UrbanCars is an artificially generated dataset with a 95% co-occurrence between car body types and relevant urban or rural backgrounds. ImageNet9 consists of 9 coarse ImageNet classes.
4.2 Model Architectures
For comparability purposes, we employ the same network architectures as those used in other works [sarridis2023flac, hong2021bb, barbano2022fairkl, sarridis2024badd]. In particular, for CelebA, we adopt the ResNet-18 architecture [he2016deep], for Waterbirds, UrbanCars, and ImageNet9 datasets, we use ResNet-50 networks. In all experiments, the projection layer is a dense layer that gets vision-language embeddings as input and its output size is aligned with the feature size of the main model.
4.3 Implementation Details
The SGD optimizer is employed for all datasets except for CelebA, where Adam optimizer is used. We use an initial learning rate of 0.001, which is divided by 10 every 1/3 of the training epochs. The weight decay is set to $10^{-4}$. The batch size is 128 for CelebA and 64 for Waterbirds, UrbanCars, and ImageNet9. Following previous works [hong2021bb, sagawa2019distributionally, li2023whac], we train the models for 40, 100, 300, and 40 for CelebA, Waterbirds, UrbanCars, and ImageNet9, respectively. For Waterbirds and UrbanCars, we do not use a learning rate scheduler. The parameters $(\alpha,\lambda)$ are (0.01, 0.5), (0.05, 0.6), (0.01, 0.4), and (0.001, 0.7) for CelebA, Waterbirds, UrbanCars, and ImageNet9, respectively. RAM [zhang2024recognize], GPT-4o [openai2023chatgpt4], OpenCLIP [clip] are employed for image tagging, irrelevant tag filtering, and vision-language encoding, respectively. For all the experiments presented in the main manuscript, we run the compared methods using the code provided by the corresponding papers, using the suggested hyperparameters. Specifically, we use 50 epochs for bias discovery and 100 upweight for JTT; $\alpha=110$ for FLAC-B; 0.1 SD coefficient; and $\alpha=0.7$ for LfF. For the baseline methods results presented in the supplementary material (i.e., closed-set), we present performance as reported in the corresponding works to avoid potential underperformance due to reimplementation. Experiments were conducted on an NVIDIA A100 GPU. All experiments were repeated for 5 different random seeds.
Table 2: Open-set performance comparison across CelebA, Waterbirds, and UrbanCars datasets.
Dataset Method WG Acc. Avg. Acc.
CelebA
LfF [nam2020LfF] 14.7 $\pm$15.2 67.1 $\pm$4.4
JTT [liu2021just] 31.5 $\pm$8.0 61.6 $\pm$8.5
SD [pezeshki2021gradient] 13.3 $\pm$8.2 67.4 $\pm$1.9
Debian [li2022discover] 12.0 $\pm$8.7 67.0 $\pm$2.6
FLAC-B [sarridis2023flac] 12.0 $\pm$8.7 65.9 $\pm$2.4
MAVias 66.7 $\pm$4.7 81.4 $\pm$1.8
Waterbirds
LfF [nam2020LfF] 30.0 $\pm$6.8 72.7 $\pm$1.4
JTT [liu2021just] 64.7 $\pm$2.4 85.2 $\pm$4.6
SD [pezeshki2021gradient] 35.0 $\pm$16.3 75.5 $\pm$4.1
Debian [li2022discover] 37.5 $\pm$0.0 74.7 $\pm$0.5
FLAC-B [sarridis2023flac] 37.5 $\pm$8.8 75.2 $\pm$2.7
MAVias 75.4 $\pm$0.9 87.5 $\pm$1.2
UrbanCars
LfF [nam2020LfF] 34.6 $\pm$2.6 61.0 $\pm$1.4
JTT [liu2021just] 69.0 $\pm$3.3 77.8 $\pm$0.3
SD [pezeshki2021gradient] 40.4 $\pm$2.7 66.5 $\pm$1.1
Debian [li2022discover] 33.2 $\pm$8.0 61.1 $\pm$2.1
FLAC-B [sarridis2023flac] 28.5 $\pm$4.3 57.3 $\pm$1.7
MAVias 84.4 $\pm$2.2 89.3 $\pm$1.3
Table 3: Open-set performance on ImageNet9 in terms of accuracy across 7 test sets.
@@ 4.3 Implementation Details
LfF [nam2020LfF] 78.70 $\pm$0.1 81.47 $\pm$0.2 61.07 $\pm$0.1 34.82 $\pm$0.2 44.46 $\pm$0.0 88.99 $\pm$0.2 94.34 $\pm$0.2
JTT [liu2021just] 84.43 $\pm$0.1 86.16 $\pm$0.5 61.09 $\pm$2.0 32.04 $\pm$1.0 36.62 $\pm$4.7 92.09 $\pm$0.5 97.71 $\pm$0.1
Debian [li2022discover] 83.02 $\pm$0.4 85.64 $\pm$0.3 64.53 $\pm$0.4 34.45 $\pm$0.1 45.00 $\pm$0.6 93.06 $\pm$0.1 97.89 $\pm$0.1
SD [pezeshki2021gradient] 87.56 $\pm$0.57 88.92 $\pm$0.74 62.60 $\pm$1.05 31.42 $\pm$2.93 40.81 $\pm$3.00 93.71 $\pm$0.71 98.16 $\pm$0.06
FLAC-B [sarridis2023flac] 84.60 $\pm$0.46 86.62 $\pm$0.45 59.84 $\pm$1.67 29.71 $\pm$0.53 40.38 $\pm$1.28 92.72 $\pm$0.73 97.89 $\pm$0.09
[…]
4.4 Evaluation Protocol
To assess bias, we primarily use worst-group accuracy (WG Acc.), which measures the accuracy of the least-performing group within a dataset, and average accuracy (Avg. Acc.), which is the mean accuracy across all groups and reflects overall model performance. The formation of these groups determines whether the evaluation supports a closed- or open-set bias scenario. Accordingly, we implement two distinct evaluation protocols. The first one is designed to align with our approach, focusing on open-set biases, while the second protocol adheres to established evaluation standards in the literature (i.e., closed-set scenario). Furthermore, since MAVias is designed for BU scenarios, we evaluate its performance against the following widely-employed and competitive BU methods: LfF [nam2020LfF], JTT [liu2021just], SD [pezeshki2021gradient], Debian [li2022discover], and FLAC-B [sarridis2023flac].
Open-set. Here, we form groups based on the presence/ absence of the detected open-set bias attributes (i.e., no predefined biases are considered). To achieve that, we use a vanilla model and the potential biases $\mathcal{B}$ extracted through MAVias to identify the subset of tags $\mathcal{B}^{\prime}\subseteq\mathcal{B}$ that actually introduce bias to the model. We define bias as occurring when the model exhibits increased accuracy on images that include a particular tag, relative to its overall accuracy across the entire dataset. After deriving $\mathcal{B}^{\prime}$ for each class within the datasets, we categorize each group of images belonging to a class into two sub-groups based on whether the samples contain at least one of the biased tags. This results in $2\times p$ groups. Subsequently, we measure the WG Acc. and Avg. Acc. across all groups. To ensure a fair assessment, the optimal training epoch for each method is selected based on the overall accuracy, without considering any information related to the biases. An exception to this protocol is ImageNet9. In the original ImageNet9 test set, we observed that for several classes, the subgroups without any biases contain very few samples (fewer than 4), making reliable evaluation difficult. We therefore use the official seven test set variations, which allow for a more comprehensive evaluation of the model’s reliance on factors beyond the target object: ORIGINAL, ONLY-BG-B (bounding box of a black object), ONLY-BG-T (bounding box of an inpainted object), NO-FG Black (segmented object removed), ONLY-FG (black background), MIXED-RAND (random background of a random class), and MIXED-NEXT (random background of the next class).
@@ 4.4 Evaluation Protocol
Closed-set. Here, we follow the protocols suggested by the dataset providers and previous works for predefined biases. In particular, for CelebA we use the accuracy of the underrepresented groups (i.e., bias-conflicting accuracy) and the average accuracy across all groups (i.e., unbiased accuracy) [sarridis2023flac, hong2021bb]. For Waterbirds, we use the WG Acc. and the Avg. Acc. across all groups. In the case of UrbanCars, we calculate the weighted average accuracy across groups, referred to as In-Distribution Accuracy (I.D. Acc), with weights determined by group representation ratios. Furthermore, I.D. Acc serves as a baseline for assessing accuracy drops related to background (BG Gap), co-occurring objects (CoObj Gap), and both background and co-occurring objects combined (BG+CoObj Gap).
@@ 4.5 Comparative Analysis
Table 2 presents the open-set performance comparison across the CelebA, Waterbirds, and UrbanCars datasets. For CelebA, MAVias significantly outperforms competing methods by achieving a 66.7% (+35.2%) WG accuracy and 81.4% (+14%) Avg. accuracy. As expected, most existing BU methods exhibit poor performance, as they struggle to handle scenarios with multiple concurrent biases [li2023whac]. For the Waterbirds, MAVias reaches a WG accuracy of 75.4% (+10.7%) and an Avg. accuracy of 87.5% (+2.3%), surpassing competing approaches by a notable margin. Similarly, in the UrbanCars dataset, MAVias achieves a WG accuracy of 84.4% (+15.4%) and an Avg. accuracy of 89.3% (+11.5%). For ImageNet9, as shown in Tab. 3, MAVias achieves accuracy improvements on the test sets with modified backgrounds, MIXED-NEXT and MIXED-RAND (+0.7% and +0.72).
Similarly, for the sets where background information is suppressed (ONLY-BG-B, ONLY-BG-T, and NO-FG) MAVias achieves improvements of +7.88%, +4.14%, and +6.82%, respectively. As observed, most competitive methods, including MAVias, underperform on ONLY-FG compared to the vanilla model. This is likely because the vanilla model can exploit biases present in the foreground (e.g., colors) to increase its accuracy. Finally, there is a 1.24% drop in performance on the ORIGINAL test set, which aligns with expectations, as MAVias is designed to rely less on shortcuts that boost overall performance. The top-10 biased tags for each Imagenet9 class are reported in Tab. 4. Moreover, Tab. 5 reports the top-10 irrelevant tags derived using MAVias for CelebA, Waterbirds, and UrbanCars datasets. For CelebA, MAVias successfully detects the primary bias (i.e., gender) while also revealing other biases associated with facial expressions (smile), shot types (selfies), and clothing items (e.g., dresses, business suits, or ties). In the case of Waterbirds, we observe that, in addition to background elements, waterbirds are predominantly depicted in flight, whereas landbirds are often observed perched on tree branches. Finally, for UrbanCars, as anticipated, the top irrelevant tags correspond to objects that are common in urban/rural environments.
Table 4: Top-10 biased tags for each class of ImageNet9.
Class Top-10 Irrelevant Tags
Bird perch, sit, branch, tree, tree branch, water, blue, twig, brown, sky
Carnivore stand, grass, stone, lush, fur, walk, lay, enclosure, tree, red
Dog brown, white, stand, black, lush, stare, green, carpet, blanket, bed
Fish sea, swim, water, underwater, aquarium, tank, coral reef, catch, man, yellow
Insect green, plant, break, floor, sit, stem, leaf, close-up, white, yellow
Instrument play, woman, ceiling, table, pillar, cloth, hang, band, tool, hand
Primate tree, sit, branch, tree branch, stare, black, enclosure, brown, stone, log
Reptile floor, stone, green, branch, tree, grass, tree branch, water, lay, sit
Vehicle road, track, red, building, curb, equipment, travel, load, train track, railroad
Table 5: Top-10 biased tags for each class of CelebA, Waterbirds, and UrbanCars.
$\mathcal{D}$ Class Top-10 Irrelevant Tags
CelebA Blonde woman, dress, pose, smile, girl, beautiful, curl, selfie, actor, carpet
Non-Blonde wear, man, shirt, selfie, tie, black, stand, stare, dress shirt, business suit
Waterbirds Landbird stand, perch, sit, tree, yellow, branch, floor, brown, forest, green
Waterbird water, stand, white, sea, fly, lake, sky, ledge, pond, stone
UrbanCars Country Car road, rural, animal, stand, white, jump, cow, lift, bull, highway
Urban Car park, house exterior, house, red, building, flip, home, lift, white, jump
For the closed-set evaluation, as reported in Tab. 6 MAVias achieves +2.7% (89.7%) Unbiased and +2.2 (87.1%) Bias-conflict accuracy compared to the second-best performing method, on CelebA. On Waterbirds, MAVias attains 93.7% WG Acc., outperforming all compared BU methods (+5%), while slightly improving the average accuracy (+0.7%). On UrbanCars, MAVias significantly enhances performance by reducing the BG, CoObj, and BG+CoObj gaps to 4.1, 2.4, and 6.7, respectively, representing absolute improvements of 4, 8.1, and 33.4. The full closed-set results are provided in the suppl. material.
Table 6: Brief closed-set performance comparison between MAVias and competitive BU methods.
Dataset Metric Best BU Method MAVias
[…]
@@
In addition, by examining the logit spaces, as shown in Fig. 3 (a), the vanilla model increases logit values proportionally with the bias, ranging from  4 to  7. This is an expected behavior, as the biases are easier to learn than the actual target and thus the model illustrates higher confidence for the bias-aligned samples. This phenomenon is referred to as Gradient Starvation [pezeshki2021gradient] and describes how bias is introduced. Also, it is theoretically shown that methodologies penalizing high logit values can mitigate this phenomenon [pezeshki2021gradient]. As shown in Fig. 3 (b), the interactions between the main model and the projection layer in the training procedure (defined by Eq. (1) and (2)) reduce the main model’s logit values proportionally to the data bias.
@@ 4.6 Ablation Analysis
An important aspect in MAVias is how the embeddings are calculated. Table 7 compares two well-known vision-language models, i.e., OpenCLIP and SigLip, and the text-based model BERT.
@@
[…]
Furthermore, Tab. 8 compares MAVias’s performance using various LLMs. GPT-4o shows the highest performance, with open-source models also performing competitively. For experiments, we use GPT-4o to minimize LLM-related errors. Table 9 presents the performance of GPT-4o in deriving relevant tags for the Waterbirds, CelebA, and UrbanCars datasets. The ground truth was manually established.
@@
[…]
Table 9: Performance of GPT-4o.
Dataset Precision Recall
Waterbirds 96.1 79.0
CelebA 81.8 75.0
UrbanCars 89.2 71.7
Furthermore, we investigate the impact of the different MAVias components. First, as discussed in Sec. 3, the logit alignment term in Eq. (2) is crucial for balancing the logits between the two models. Table 10 shows the impact of removing this term from the training on the effectiveness of MAVias.